# Causality for Earth Science - A Review on Time-series and Spatiotemporal Causality Methods

### v1

@@ 3.5.1 Mapped-PCMCI
(Tibau et al., 2022) presented a spatiotemporal stochastic climate model SAVAR which can be used to benchmark causal discovery methods for teleconnections, providing insights into the strengths and weaknesses of different analysis methods. The authors also introduced novel causal discovery method named Mapped-PCMCI that outperforms existing approaches, contributing to the advancement of process-based understanding and climate model evaluation. The proposed SAVAR model benchmarks different teleconnection analysis methods including grid-level causal discovery methods and a combination of dimension-reduction and causal discovery methods. The experiments conducted in the paper demonstrate that the grid-level causal method based on the SAVAR model outperforms baseline causal discovery methods, which do not consider the mode structure and attempt to directly infer the causal graph at the grid level. Proposed with the SAVAR is the Mapped-PCMCI method, which is based on the assumption that the causal dependencies within a gridded dataset have a lower-dimensional mode representation. The method consists of four steps (Figure 5): (1) Perform a dimensionality reduction method on the gridded data to extract a limited number of mode time series variables. (2) Apply a causal discovery method to the mode time series variables to obtain the parents and the estimated causal network of the modes. (3) Estimate (lagged) causal effects for all links to obtain a coefficient matrix. (4) Invert the dimension reduction by mapping the causal effects among the grid locations using the modes’ weights. This method aims to overcome the challenges of dealing with large networks, nonstationary networks, and the high-dimensional and redundant estimation problem at the grid level. It provides a spatial grid-level representation of the causal network obtained at the mode level, allowing for the analysis of causal relationships in complex systems. Python code for both the SAVAR and Mapped-PCMCI were published on GitHub  [^8] https://github.com/xtibau/.
@@ 3.5.2 Interactive Causal Structure Discovery (ICSD)
(Melkas et al., 2021) presented a workflow ICSD (Interactive Causal Structure Discovery) in Earth System Sciences that takes into account experts’ domain knowledge during the application of causal discovery (CD) algorithms in Earth sciences systems. ICSD provides users an interactive way to edit the outputs of different CD algorithms and iteratively incorporate prior knowledge in the initial output models. It formulates the structure learning problem as building a probabilistic model of the data and encodes expert’s prior knowledge as a prior distribution over all possible causal structures. The current formulation takes a greedy approach which leads to local optima where the choice of the initial state affects the final model. A score is associated with each causal model that reflects the log-likelihood of the model. For demonstration of the approach, three different CD algorithms (PC, GES and LiNGAM) were experimented on a forestry dataset that includes daytime measurements of variables such as shortwave downward radiation, temperature, latent heat flux vapor, pressure deficit, etc. Experimental results suggest that even a small amount of prior knowledge is useful in improving the results of CD algorithms. The experiments also showed how overfitting and concept drift can occur and be detected by ICSD. At present, the user navigation begins with the outputs from different CD algorithms which is sub-optimal. A future work is to find an optimal set of representative starting points for exploration.
@@ 3.5.3 Spatio-Temporal Causal Discovery Framework(STCD)
The Spatio-Temporal Causal Discovery Framework (STCD) (Sheth et al., 2022) aimed to identify causal relationships in the spatio-temporal domain by enforcing both temporal and spatial constraints. This framework is motivated by the influence of causal factors in hydrological systems where the geographical location of a river or a subsidiary is a crucial factor in deciding the causal parents. STCD primarily extends the TCDF (Nauta et al., 2019b) framework by adding a component that ensures the enforcement of the spatial constraint. TCDF is a temporal causal discovery approach based on CNNs for multivariate time series data that uses an attention mechanism to identify potential causal parents of the target time-series. Using the attention scores, it filters the list of potential candidates. The proposed framework STCD on the other hand eliminates irrelevant candidates by penalizing the attention scores if the candidates violate the spatial constraints. When the attention vector for a target time series is obtained, then the spatial constraint is imposed on it. In hydrological systems, a river located geographically below the target river can never be a causal parent even if they satisfy the temporal constraint because the flow of water is always from top to bottom. Motivated by this idea, the spatial constraint imposes a direction of the flow. Specifically, it is the product of a distance $d$ and a spatial coefficient $\lambda$. The distance has either a positive or negative direction based on the geographical positioning of the locations, and $\lambda$ controls the effect of the spatial constraint on the attention score. The framework is tested to discover causal relationships of streamflow at the mouth of the Brazos basin in Texas from runoff in the basin. A limitation of STCD is that it models the spatial and temporal interactions separately which might hinder the discovery process when the spatial and temporal components influence each other.
@@ 3.5.4 Group Elastic Net
(Lozano et al., 2009) proposed a data-centric approach to climate change attribution, using spatial-temporal causal modeling to analyze climate observations and forcing factors. The authors developed a novel method called Group Elastic Net to infer causality from the data and incorporate extreme value modeling to study extreme climate events. With the assumption of spatial stationarity, the model combines graphical modeling techniques with Granger Causality to derive effective methods for causal modeling based on the spatio-temporal structure of the data and enforces sparsity at the group level. The model also applies extreme-value theory to model extreme events and incorporates these estimates into the causal modeling and attribution process. The experiments involve data collection from multiple sources. For real data collection, the researchers compiled a comprehensive set of relevant variables for climate modeling in North America. They obtained data from various sources including the Climate Research Unit (CRU), the National Oceanic and Atmospheric Administration (NOAA), NASA, the National Climate Data Center (NCDC), and the Carbon Dioxide Information Analysis Center (CDIAC). These sources provided data on climate variables such as temperature, precipitation, solar irradiance, greenhouse gases, and aerosols. In addition to the real climate data, the researchers also conducted simulation experiments using synthetic spatial-temporal data. They used a spatial-temporal vector autoregressive (VAR) model to generate the synthetic data. The experiments compared the performance of the Group Elastic Net method, which considers spatial interactions, with a method that neglects such interactions. The experimental results indicate that changes in temperature are not solely accounted for by solar radiance but are attributed more significantly to CO2 and other greenhouse gases. The results also show a significant increase in the intensity of extreme temperatures, and these changes are attributable to greenhouse gases. The proposed approach offers a useful alternative to simulation-based climate modeling and provides valuable insights from a fresh perspective.
@@
Previous methods for identifying causal pathways in air pollution have been proposed from two perspectives: pattern-based and Bayesian-based. Pattern-based approaches provide shallow understanding and have limited usability due to a large number of patterns, while Bayesian-based approaches are limited by noise, data sparsity, computational cost, and confounding factors. Zhu et al. (Zhu et al., 2017) present a novel approach, called pg-Causality, which combines pattern mining and Bayesian learning to efficiently identify spatiotemporal causal pathways for air pollutants using urban big data. The approach overcomes the challenges of noise, computational complexity, and complex causal pathways, and thus outperforming traditional methods in terms of time efficiency, inference accuracy, and interpretability. The authors use the FEP Mining Algorithm to mine frequent episode patterns (FEPs) in a symbolic pollution database. This algorithm considers constraints such as consecutive symbols being different and a specified temporal constraint between consecutive records. After discovering the FEPs, the authors extract candidate causes for each sensor by finding pattern-matched pairs within a specified time lag threshold. The authors use a Gaussian Bayesian network (GBN) based graphical model to capture the causal relationships among air pollutants. They generate initial causal pathways by incorporating the extracted matched patterns and candidate sensors into the GBN model. Finally, the authors refine the causal structures using an EM learning phase and a structure reconstruction phase. The EM learning phase learns the parameters of the graphical model, while the structure reconstruction phase selects the top neighborhood sensors based on the newly generated GCscore and updates the Q matrix. The experiments were conducted using real-world data sets from North China, Yangtze River Delta, and Pearl River Delta. These data sets included records of 6 air pollutants and 5 meteorological measurements. The approach showed scalability in identifying causal pathways for air pollutants at the sensor level, which is more than ten times larger than at the city level. The overall experimental results showed that the proposed approach achieved high accuracy, time efficiency, and scalability in inferring causal relationships in air pollution data.
@@ 3.6.1 Applications of Granger Causality
In climate research, understanding complex phenomena, such as teleconnection patterns, is important because it links atmospheric changes in one region to impacts in distant regions. However, the automatic identification of these patterns from observational data is still unresolved due to nonlinearities, nonstationarities, and the limitations of correlation analyses. Varando et al.(Varando et al., 2021) propose a deep learning approach to address these problems and learn Granger causal feature representations that capture the true causal effects of the target index, such as El Niño Southern Oscillation (ENSO) or North Atlantic Oscillation (NAO). The authors propose a method called the Granger Penalized Autoencoder with the assumptions including the presence of nonlinearities and nonstationarities in the observational data, as well as the limitation of correlation analyses in identifying true causal patterns. By utilizing normalized difference vegetation index (NDVI) data collected from MODIS reflectance data over 11 years, their work identified clear patterns of the causal footprints of ENSO on vegetation in different regions. The GitHub repository of this research is approachable [^9] https://github.com/IPL-UV/LatentGranger. (Mosedale et al., 2006) used a Granger causality based approach to quantitatively measure the feedback of daily sea surface temperatures (SSTs) on daily values of the North Atlantic Oscillation (NAO). This was done by simulating a realistic coupled general circulation model (GCM). This study is an extension of the work by Mosedale et al. in 2005. where the Granger causality approach is used to find the best time series models for modeling the coupled system for greater flexibility. (Smirnov and Mokhov, 2009) introduced the idea of long-term causality, which is an extension of Granger causality. Long-term causality was estimated from data through empirical modeling and analysis of model dynamics under different conditions. They applied this concept to find out how strongly the global surface temperature (GST) is affected by variations in carbon dioxide, atmospheric content, solar activity, and volcanic activity during the last 150 years. (Kodra et al., 2011) extended the classic Granger causality test to handle the multisource nature of data. Using a reverse cumulative Granger causality test, they tested the hypothesis that Granger causality can be extracted from the bivariate series of globally averaged land surface temperature (GT) observations and observed CO2 in the atmosphere.
@@ 3.6.2 Applications of Pearl Causality
Causal discovery algorithms based on probabilistic graphical models have been applied in geoscience applications to identify and visualize dynamical processes, but the lack of ground truth and unexplained connections have posed several challenges. To address these challenges, Ebert-Uphoff et al.(Ebert-Uphoff and Deng, 2017) developed a simulation framework using synthetic spatio-temporal data to better understand the physical processes and interpret the resulting connectivity graphs, ultimately solving the mystery of the previously unexplained connections. This approach allows for the resolution of previously unexplained connections and provides a benchmark for other causal discovery algorithms. The authors used a constraint-based structure learning method called the PC stable algorithm, which is a modification of the classic PC algorithm. This algorithm produces graph structures where observed variables form the nodes, and connections between nodes indicate potential cause-effect relationships. The PC stable algorithm has advantages such as increased robustness of results and suitability for parallelization. The authors dropped the requirement of causal sufficiency and focused on necessary conditions for cause-effect relationships. The article discusses the results of three different scenarios in the simulations. In Scenario 1, increasing the spatial resolution leads to more edges for a specific time interval and fewer edges for a longer time interval. In Scenario 2, concurrent edges are believed to be caused by contradictory velocities at the boundaries of the advection field. In Scenario 3, concurrent edges in the center do not match the typical diffusion pattern and appear to fill modeling gaps. The velocity estimates in Scenario 1 with higher speed are weak due to many high-speed interactions represented as concurrent edges.
@@ 4 Causal Inference
Causal inference can be defined as the process of estimating the causal effects (influence) of one event, process, state or object (a cause) on the another event, process, state or object (an effect). Causal inference has been applied to study environmental science for several decades, with early applications dating back to the mid-20th century (Hill, 1965). However, much of the earlier causal inference based analysis was done on independent and identically distributed (i.i.d) data utilizing statistical inferential and regression techniques to estimate the causal effects on potential outcomes (Pearl, 2009). This section presents the key terminologies (Table 7), common approaches and causal assumptions required to perform causal inference techniques. We explain the different time-series and spatiotemporal causal inference techniques introduced, along with their limitations and applications in Earth Science.
@@ 4.1 Common Approaches
For estimation of causal effect, there are two main categories of techniques, potential outcome framework and do-calculus.
4.1.1 Potential Outcome Framework:
The potential outcome framework relies on hypothetical interventions such that it defines the causal effect as the difference between the outcomes that would be observed with and without exposure to the intervention. This technique is widely used in epidemiology where patients are randomly divided into treated and controlled groups and the effectiveness of treatment is inferred by observing patients condition with and without undergoing a treatment (Rubin, 2005). The treatment effect can be measured at individual, treated group, sub-treated group and entire population levels.
@@ 4.2 Assumptions for Causal Inference
Under the consistency condition, the potential outcome for the treated subject $Y_{X=1}$ is considered equal to the observed outcome Y. The same goes for the untreated subject.
@@
Under the conditional exchangeability assumption, also known as ”weak ignorability”, the condtional probability of receiving treatment depends only on the covariates $Z$, that is, $Y_{x}$ and treatment $X$ are are statistically independent given every possible value of $Z$. On the contrary, unconditional exchangeability implies that treatment group has the same distribution of outcomes as the untreated control group.
@@ 4.3.1 Root Mean Square Error (RMSE)
Researchers evaluate the performance of their predictive models using the Root Mean Square Error (RMSE) which can be only calculated for factual observational data but cannot be done for counterfactual predictions. Since the ground truth information is only available for synthetic data, we further evaluate the causal effect estimation skill of a model using the PEHE metric.
@@ 4.3.2 Precision in Estimated Heterogeneous Effects (PEHE)
This metric is commonly used in machine learning literature for calculating the average error across the predicted average treatment effects (ATEs) (Hill, 2011).
@@ 4.4 Time-series Causal Inference
Majority of the causal inference models work in time and space invariant settings. However, when it comes to time-series data, the question changes to inferring the effect of treatment $X$ on outcome $Y$ at time $t$ in the presence of a set of covariates $Z$. Such settings require models that can estimate time-varying causal or treatment effects. This further leads to the problem of time-varying confounding, that is the common influence a past treatment $X_{t}$ or covariate $Z_{t}$ might have on the future treatments $X_{t+1}$ and the future outcome $Y_{t+1}$. Traditional methods for performing time-series causal inference typically involve statistical modeling techniques that aim to identify causal relationships between variables in time-series data, however, these methods have several limitations, which can affect the reliability and validity of causal conclusions drawn from the analysis.
In causal inference, confoundedness poses a significant challenge because it can lead to incorrect conclusions about the causal relationship between the independent variable and the dependent variable. When confounding variables are not properly accounted for, the observed association between the independent and dependent variables may be due to the confounding variable rather than a true causal effect. Balancing scores that incorporate propensity score are the most common approaches to debias confounding effects. To overcome this challenge, Robin’s g-methods have shown to provide promising results on reducing bias caused by time-varying treatment and covariates on the potential outcome (Naimi et al., 2017). The prediction models of these estimators are typically based on linear or logistic regression, the downside of which is that in case of complex non-linearities in treatment or outcome variables, these methods will lead to inaccurate results. G-methods provide metrics to overcome the problem of time-varying confounding through standardization, g-computation and inverse probability of treatment weighted (IPTW) estimators (Naimi et al., 2017).
4.5 Time Series Causal Inference Methods
So far, we have seen causal inference surveys focusing on CI methods for i.i.d. data (Table 1), however, real world observations require dynamic or time-varying causal analysis focusing on calculating the impact of an intervention on a sequential or time-varying outcome (Moraffah et al., 2021). For instance, policymakers and climate change activists would be interested in identifying the impact of lowering CO2 emissions on the rate of ozone depletion over a specific period of time. Depending upon the underlying methodology, time-series causal inference methods can be categorized into (i) time-varying causal inference, and (ii) time-invariant causal inference methods. We summarize some of the widely-used causal inference methods under each category below.

### v2

@@ 3.5.1 Mapped-PCMCI
(Tibau et al., 2022) presented a spatiotemporal stochastic climate model SAVAR which can be used to benchmark causal discovery methods for teleconnections, providing insights into the strengths and weaknesses of different analysis methods. The authors also introduced novel causal discovery method named Mapped-PCMCI contributing to causality-based climate model evaluation. Mapped-PCMCI is based on the assumption that the causal dependencies within a gridded dataset can have a lower-dimensional latent representation. The method consists of four steps (Figure 5): (i) Dimensionality reduction method on the gridded data, (ii) Application of causal discovery method to the lower-dimensional time series variables, (iii) Estimation of (lagged) causal effects, (iv) Inverse transformation of lower-dimension features back to the original grid locations. This method aims to overcome the challenges of dealing with large networks, nonstationary networks, and the high-dimensional causal estimation of spatiotemporal data. It provides a spatial grid-level, allowing for the analysis of causal relationships in complex systems. Python code for both the SAVAR and Mapped-PCMCI is accessible on GitHub  [^8] https://github.com/xtibau/.
@@ 3.5.2 Interactive Causal Structure Discovery (ICSD)
Melkas et al. (2021) introduced a method called Interactive Causal Structure Discovery (ICSD) designed to integrate expert knowledge during the application of causal discovery algorithms in Earth science. ICSD allows users to interactively modify initial algorithm outputs, enabling the iterative inclusion of expert knowledge into initial models. This method treats structure learning as a task of creating a probabilistic model based on the data, incorporating expert knowledge as prior distributions across possible causal structures. The current method uses a greedy approach, which can get stuck in local optima due to the dependence on the initial model state. Each causal model is evaluated based on a score reflecting its log-likelihood. For practical illustration, ICSD was applied to a forestry dataset with variables like shortwave downward radiation and temperature using three causal discovery algorithms: PC, GES, and LiNGAM. The findings indicate that even minimal prior knowledge can enhance the performance of these algorithms, and highlighted the potential for overfitting and concept drift, which ICSD can identify and address. Currently, the system starts with outputs from various causal discovery algorithms, which is not ideal. Future work aims to identify optimal starting points for user exploration.
@@ 3.5.3 Spatio-Temporal Causal Discovery Framework(STCD)
The Spatio-Temporal Causal Discovery Framework (STCD) (Sheth et al., 2022) aims to identify causal relationships in spatiotemporal data by enforcing both temporal and spatial constraints. STCD primarily extends the TCDF (Nauta et al., 2019b) framework by adding a component that ensures the enforcement of the spatial constraint. Using the TCDF attention scores, it filters the list of potential candidates. The proposed framework STCD eliminates irrelevant candidates by penalizing the attention scores if the candidates violate the spatial constraints. This framework takes influence from hydrological systems, where a river’s geographic location gives it temporal precedence over other rivers, based on the direction of water flow from one river to the other. Motivated by this idea, the spatial constraint imposes a direction of the flow. Specifically, it is the product of a distance $d$ and a spatial coefficient $\lambda$. The distance has either a positive or negative direction based on the geographical positioning of the locations, and $\lambda$ controls the effect of the spatial constraint on the attention score. A limitation of STCD is that it cannot model the spatial and temporal interactions simultaneously, which might hinder the discovery process when the spatial and temporal components influence each other.
@@ 3.5.4 Group Elastic Net
(Lozano et al., 2009) proposed a data-centric approach to climate change attribution, using spatial-temporal causal modeling to analyze climate observations and forcing factors. The authors developed a novel method called Group Elastic Net to infer causality from the data and incorporate extreme value modeling to study extreme climate events. With the assumption of spatial stationarity, the model combines graphical modeling techniques with Granger Causality for spatiotemporal data. The model also applies extreme-value theory to model extreme events and incorporates these estimates into the causal modeling and attribution process. The experiments involve data collection from multiple sources. For real-world data collection, the researchers obtained data from various sources including the Climate Research Unit (CRU), the National Oceanic and Atmospheric Administration (NOAA), NASA, the National Climate Data Center (NCDC), and the Carbon Dioxide Information Analysis Center (CDIAC). These sources provided data on climate variables such as temperature, precipitation, solar irradiance, greenhouse gases, and aerosols. In addition to the real climate data, the researchers used a spatial-temporal vector autoregressive (VAR) model to generate synthetic data. The experiments compared the performance of the Group Elastic Net method, which considers spatial interactions, with a method that neglects such interactions. Utilizing the method, the researchers estimated the influence of CO2 and greenhouse gases on increasing temperatures. Their proposed approach can offer a useful alternative to simulation-based climate modeling.
@@
Zhu et al. (Zhu et al., 2017) presented a novel approach, called pg-Causality, to efficiently identify spatiotemporal causal pathways for air pollutants using urban big data by combining pattern mining and Bayesian learning. The approach overcomes the challenges of noise, computational complexity, and complex causal pathways, and thus outperforming traditional methods in terms of time efficiency, inference accuracy, and interpretability. The authors use the FEP Mining Algorithm to mine frequent episode patterns (FEPs) in a symbolic pollution database. This algorithm considers constraints such as consecutive symbols being different and a specified temporal constraint between consecutive records. After discovering the FEPs, the authors extract candidate causes for each sensor by finding pattern-matched pairs within a specified time lag threshold. The authors use a Gaussian Bayesian network (GBN) based graphical model to identify causal pathways among air pollutants. The authors refine the causal structures learnt from GBN using an expectation-maximization (EM) learning phase and a structure reconstruction phase. The experiments were conducted using real-world dataset of 6 air pollutants and 5 meteorological measurements from from North China, Yangtze River Delta, and Pearl River Delta. The experimental results demonstrated pg-Causality’s high accuracy, time efficiency, and scalability in inferring causal relationships in air pollution data.
@@ 3.6.1 Applications of Granger Causality
In climate research, understanding complex phenomena, such as teleconnection patterns, is important because it links atmospheric changes in one region to impacts in distant regions. However, the automatic identification of these patterns from observational data is still unresolved due to nonlinearities, nonstationarities, and the limitations of correlation analyses. Varando et al.(Varando et al., 2021) propose a deep learning approach to address these problems and learn Granger causal feature representations that capture the true causal effects of the target index, such as El Niño Southern Oscillation (ENSO) or North Atlantic Oscillation (NAO). The authors propose a method called the Granger Penalized Autoencoder with the assumptions including the presence of nonlinearities and nonstationarities in the observational data, as well as the limitation of correlation analyses in identifying true causal patterns. By utilizing normalized difference vegetation index (NDVI) data collected from MODIS reflectance data over 11 years, their work identified clear patterns of the causal footprints of ENSO on vegetation in different regions. The GitHub repository of this research is approachable [^9] https://github.com/IPL-UV/LatentGranger. (Mosedale et al., 2006) used a Granger causality based approach to quantitatively measure the feedback of daily sea surface temperatures (SSTs) on daily values of the North Atlantic Oscillation (NAO). This was done by simulating a realistic coupled general circulation model (GCM). This study is an extension of the work by Mosedale et al. in 2005. where the Granger causality approach is used to find the best time series models for modeling the coupled system for greater flexibility. (Smirnov and Mokhov, 2009) introduced the idea of long-term Granger causality to find out how strongly the global surface temperature (GST) is affected by variations in carbon dioxide, atmospheric content, solar activity, and volcanic activity during the last 150 years. (Kodra et al., 2011) extended the classic Granger causality test and hypothesized the multisource causal relation between globally averaged land surface temperature (GT) observations and observed CO2 in the atmosphere.
@@ 3.6.2 Applications of Pearl Causality
Causal discovery algorithms based on probabilistic graphical models have been applied in geoscience applications to identify and visualize dynamical processes, but the lack of ground truth and unexplained connections have posed several challenges. To address these challenges, Ebert-Uphoff et al.(Ebert-Uphoff and Deng, 2017) developed a simulation framework using synthetic spatiotemporal data to better understand the physical processes and interpret the resulting connectivity graphs, ultimately solving the mystery of the previously unexplained connections. This approach allows for the resolution of previously unexplained connections and provides a benchmark for other causal discovery algorithms. The authors used a constraint-based structure learning method called the PC stable algorithm, which is a modification of the classic PC algorithm. The PC stable algorithm has advantages such as increased robustness of results and suitability for parallelization. The authors dropped the requirement of causal sufficiency and focused on necessary conditions for cause-effect relationships. The article discusses the results of three different scenarios in the simulations. In Scenario 1, increasing the spatial resolution leads to more edges for a specific time interval and fewer edges for a longer time interval. In Scenario 2, concurrent edges are believed to be caused by contradictory velocities at the boundaries of the advection field. In Scenario 3, concurrent edges in the center do not match the typical diffusion pattern and appear to fill modeling gaps. The velocity estimates in Scenario 1 with higher speed are weak due to many high-speed interactions represented as concurrent edges.
@@ 4 Causal Inference
Causal inference or causal effect estimation is the process of quantifying the causal impact of an entity (state or event) that contributes to the data-generation process of another entity (outcome, state or event). Causal inference has been applied to study environmental science for several decades, with early applications dating back to the mid-20th century (Hill, 1965). However, much of the earlier causal inference based analysis was done on independent and identically distributed (i.i.d) data utilizing statistical inferential and regression techniques to estimate the causal effects on potential outcomes (Pearl, 2009). This section presents the key terminologies (Table 7), common approaches and causal assumptions required to perform causal inference techniques. We explain the different time-series and spatiotemporal causal inference techniques introduced, along with their limitations and applications in Earth science.
@@ 4.1 Common Approaches
Causal inference methods are based on one of the two main categories: do-calculus and potential outcome framework.
@@ 4.1.1 Do-Calculus
4.1.2 Potential Outcome Framework:
The potential outcome framework, also known as the Rubin Causal Model (RCM), is a formal approach to causal inference in statistics and social sciences (Rubin, 2005). It provides a way to define and estimate causal effects by considering hypothetical scenarios, also known as treatments. For each unit (e.g., individual, event, etc.) in the study, the potential outcome framework considers the outcomes under each possible treatment. If there are two treatments, $T=1$ and $T=0$, then each unit can have one of two potential outcomes: $Y(1)$: the outcome of the unit after undergoing treatment, and $Y(0)$: the outcome of the unit without receiving treatment. For any unit, we can only observe one of the potential outcomes (the one corresponding to the treatment actually received). The other potential outcome is counterfactual and unobservable. This makes direct calculation of causal effects impossible without further assumptions or methods.
@@ 4.2 Assumptions for Causal Inference
Under the consistency condition, the potential outcome for the treated subject $Y_{X=1}$ is considered equivalent to the observed outcome Y. The same goes for the untreated subject.
@@
Under the conditional exchangeability assumption, also known as ”weak ignorability”, the condtional probability of receiving treatment depends only on the covariates $Z$, that is, $Y_{x}$ and treatment $X$ are statistically independent given every possible value of $Z$. On the contrary, unconditional exchangeability implies that treatment group has the same distribution of outcomes as the untreated control group.
@@ 4.3.1 Root Mean Square Error (RMSE)
The RMSE score is a widely used evaluation metric to calculate error in regression-based predictions. In case of causal inference, this metric is used for evaluating the performance of predictive models when ground truth information about factual and counterfactual data is known. RMSE error can be computed using the formula in Equation  5 where $Y_{i}$ represents the actual value of the outcome and $\hat{Y_{i}}$ represents the estimated value of the $i^{th}$ potential outcome by a causal inference method.
@@ 4.3.2 Precision in Estimated Heterogeneous Effects (PEHE)
PEHE is a metric similar to the RMSE score that is used for calculating the error in the actual and predicted average treatment effects (ATEs) (Hill, 2011). Just like RMSE score, PEHE scores are only applicable when counterfactual values are already known, a case common in synthetic datasets.
@@ 4.4 Time-series Causal Inference
Real world observations require dynamic or time-varying causal analysis focusing on calculating the impact of an intervention on a sequential or time-varying outcome (Moraffah et al., 2021). For instance, policymakers and climate change activists would be interested in identifying the impact of lowering CO2 emissions on the rate of ozone depletion over a specific period of time. Time series causal inference provides a framework for understanding and quantifying causal relationships in data that evolve over time, employing a range of statistical and econometric techniques to address the unique challenges posed by temporal data such as non-stationarity and time-varying confoundedness. A non-stationary time series is a time series whose statistical properties, such as mean, variance, and autocorrelation, change over time. Non-stationarity exhibits spurious correlations in data that leads to incorrect causal effect estimations. Confoundedness also comes as a significant challenge because it can lead to incorrect conclusions about the causal relationship. Confounders are variables (or covariates) whose past values influence the future values of both the cause (treatment) and the effect (outcomes). When confounding is not properly accounted for, the observed association between the cause and effect may be due to the confounding variable rather than a true causal effect leading to biased predictions. Both non-stationarity and confounding are prevailing challenges in causal inference for Earth science applications.
We summarize some of the widely-used causal inference techniques for time-varying and time-invariant data, and their applications in Earth science, below.